find the direction in which a straight line must be drawn through the point (1,2),so that its point of intersection with the line x+y=4 is at a distance ^2/3 from the given point.

Grade:11

1 Answers

AAYUSH KESHAVA
11 years ago

let l1 be req line and l2 be the given one

since l1 itersects l2 , let the coordinates of intersection be (x,4-x)

A/Q

[(x-4)-2]2 + [x-1]2 =(2/3)2

=> x2+36+12x+x2+1-2x=4/9

now solve the eqn and hunt for answer!!!!!!!!